OpenStudy (anonymous):

If you are trying to find a power series representation of a function f(x) and f'(x) is a known series, x^n then the power series rep of f(x) is the derivative of the series x^n, correct?

OpenStudy (anonymous):

f'(x) - x^n f(x)- 1/n+1 x^(n+1)

OpenStudy (anonymous):

Yes, for only x^n other cases are different. P.S. Remember to differentiate right. f(x) = x^n, f'(x) = nx^(n-1)

OpenStudy (anonymous):

see thats what I got pk, but the answer my prof gave is what imran showed.. which would be the integral?

OpenStudy (anonymous):

to go from f'(x) to f(x) you have to integrate , reverse of f(x) to f'(x) where you differentiate

OpenStudy (anonymous):

He is right, I misread your question.

OpenStudy (anonymous):

oh so if I have to differentiate f(x) to get a known power series then i take the integral of that power series

OpenStudy (anonymous):

and if i have to take the integral of f(x) to get a known power series then i differentiate the series?

OpenStudy (anonymous):

well if they give you f'(x) and they want f(x) you would integrate

OpenStudy (anonymous):

well the problem i was given was \[\ln (1+x^4)\] and I took the derivative of this.. and then found then differentiated the power series.. But now that I am looking at it I could have just integrated the known power series for 1/1-x and that would have gave me the power series for ln(1-x) and then i just plug in x^4; but I would assume both ways would have to give you equivalent series?

OpenStudy (anonymous):

well id have to plug in -x^4

OpenStudy (anonymous):

-ln(1-x)= 1/n+1 x^(n+1)

OpenStudy (anonymous):

ln(1-x)=-1/(n+1) x^(n+1)

OpenStudy (anonymous):

replace x with -x^4

OpenStudy (anonymous):

\[\frac{-1}{n+1}\left(-x^4\right)^{n+1}\]

OpenStudy (anonymous):

\[\sum _{n=0}^{\infty } \frac{-1}{n+1}\left(-x^4\right)^{n+1}\]

OpenStudy (anonymous):

thanks, I see it now... sorta lol